OpenStudy (anonymous):

PLEASE ANYONE I REALLY NEED TO GET DONE there will definitely be a reward for helping!

OpenStudy (anonymous):

[9.05] How many and of what type are the solutions to x2 + 10x + 25 = 0? No real solutions Two identical rational solutions Two different rational solutions Two irrational solutions 15. [9.05] What are the exact solutions of x2 + 5x + 2 = 0? x = the quantity of negative 5 plus or minus the square root of 33 all over 2 x = the quantity of 5 plus or minus the square root of 33 all over 2 x = the quantity of negative 5 plus or minus the square root of 17 all over 2 x = the quantity of 5 plus or minus the square root of 17 all over 2 16. [9.05] Solve 2x2 + 5x + 5 = 0. Round solutions to the nearest hundredth. No real solutions x ≈ −3.27 and x ≈ 0.77 x ≈ −2.98 and x ≈ −7.02 x ≈ −0.77 and x ≈ 3.27 17. [9.05] What are the approximate solutions of 3x2 − 5x = 3 rounded to the nearest hundredth? No real solutions x ≈ −1.41 and x ≈ 6.41 x ≈ −0.47 and x ≈ 2.14 x ≈ −2.14 and x ≈ 0.47 18. [9.06] Katie wants to hang a painting in a gallery but the painting and frame must have an area of 45 square feet. The painting is 6 feet wide by 7 feet long. The frame must have an equal width of x on each side. Which quadratic equation represents the area of the painting and frame combined? 4x2 + 26x + 45 = 0 4x2 + 26x − 3 = 0 x2 + 13x − 3 = 0 x2 + 13x + 45 = 0 19. [9.06] A firecracker shoots up from a hill 160 feet high with an initial speed of 105 feet per second. Using the formula H(t) = −16t2 + vt + s, determine how long it will take the firecracker to hit the ground. 7.5 seconds 7.8 seconds 8.1 seconds 8.3 seconds 20. [9.06] Derek kicks a soccer ball off the ground and in the air with an initial velocity of 31 feet per second. Using the formula H(t) = −16t2 + vt + s, what is the maximum height the soccer ball reaches? 14.2 feet 14.6 feet 15.0 feet 15.3 feet 21. [9.06] Using graphing technology, approximate the solutions of 0.55x2 − 0.33x − 5.6 = 0. x ≈ −2.91 and x ≈ 3.51 x ≈ −3.14 and x ≈ 3.81 x ≈ −2.76 and x ≈ 3.62 x ≈ −2.61 and x ≈ 3.84 22. [9.06] Using graphing technology, approximate where the graph of y = 5x2 − 3x − 9 crosses the x-axis. (1.55, 0) and (−1.15, 0) (1.23, 0) and (−1.26, 0) (1.32, 0) and (−1.01, 0) (1.68, 0) and (−1.08, 0)